OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ±5/6 x.

OpenStudy (agent0smith):

you'll want to use this again http://www.jfkcougars.org/ourpages/auto/2013/2/5/43231143/Conics-Formula-Sheet.pdf since the vertices are (0, 10) and (0, -10) its a vertical hyperbola. Center is (0, 0) since that's halfway between the vertices.

OpenStudy (anonymous):

(y/b)^2-(x/a)^2 = 1 (y/10)^2-(x/a)^2= 1 lim(f(x)/x)=m x-infinity (y/10)^2- (x/a)^2= 1 y = square root of (100 + 100(x/a)^2) lim (suare root(100 + 100(x/a)^2)/x ) =10/a x-infinity 10/a =5/6 a =12 (y/10)^2- (x/12)^2= 1 @agent0smith

OpenStudy (agent0smith):

That looks right to me